ToxicaDennisIce:

Martha states that –6 is a rational number. Which is a correct explanation for this statement? The number –6 is the opposite of 6. The number –6 is a negative integer. The number –6 can be written as StartFraction negative 6 over 1 EndFraction. The number –6 is less than 0.

jimthompson5910:

hint: all rational numbers can be written in the form of p/q p and q are placeholders for integers. The value of q cannot be 0 For example, 2/3 is a rational number since p = 2 and q = 3

ToxicaDennisIce:

jimthompson5910:

As another example, lets say p = 7 and q = 10 This means \(\large\frac{7}{10}\) is a rational number since we started with \(\large\frac{p}{q}\) and replaced those letters with the mentioned values.

ToxicaDennisIce:

i dont get it

jimthompson5910:

Think of any fraction and tell me what you're thinking

ToxicaDennisIce:

The number –6 can be written as StartFraction negative 6 over 1 EndFraction.

ToxicaDennisIce:

this???

jimthompson5910:

yes because \(\large -6=\frac{-6}{1}\) we have shown that -6 is a rational number. So Martha is correct

jimthompson5910:

If the term "rational number" seems confusing, just think "fraction" and it might help

ToxicaDennisIce:

yea thx

jimthompson5910:

no problem